OpenStudy (anonymous):

A line is perpendicular to x+2y+2z=0 and passes through (0,1,0) . The perpendicular distance of the line from the origin is ?

OpenStudy (amistre64):

oh the suspense ....

OpenStudy (amistre64):

or is that the question .....

OpenStudy (anonymous):

That is the complete question !

OpenStudy (anonymous):

Better :D

OpenStudy (amistre64):

pfft, now the suspense is gone lol

OpenStudy (amistre64):

do you know what line is always perp to a plane?

OpenStudy (amistre64):

by line i mean vector ... we will define the line using the vector and the stated point

OpenStudy (anonymous):

The coefficients of x,y,z gives the direction ratio of the normal

OpenStudy (anonymous):

Ah. :| Sorry , please continue .

OpenStudy (amistre64):

fine then, ill just take my brillant ideas and go make a sammich .... oh you still need me then eh :)

OpenStudy (amistre64):

yes, the coeffs now we have a vector, we want another vector, specifically the one for the stated point to the origin

OpenStudy (anonymous):

The equation of the line is (0,1,0) +k(1,2,2)

OpenStudy (amistre64):



OpenStudy (amistre64):

how do we determine the angle between 2 vectors?

OpenStudy (anonymous):

Dot product..

OpenStudy (amistre64):

there is a dotproduct equivalnce yes, namely: \[|u|~|v|~cos(\alpha)=u\cdot v\]

OpenStudy (amistre64):

sin(a) = d/h d = h sin(a)

OpenStudy (anonymous):

Yep ! Agreed.

OpenStudy (anonymous):

Thanks a lot :D

OpenStudy (amistre64):

so vector form (0,1,0) to (0,0,0) = (0,-1,0) of length 1 vector of normal is (1,2,2) of length sqrt(9)=3 cos(a) = (0,-1,0).(1,2,2)/3 a = cos^(-1) (-2/3) sin(a) = sin(cos^(-1) (-2/3)) and since our to the origin is a length of 1 it really not that important. d = sin(cos^(-1)(-2/3))

OpenStudy (amistre64):

good luck ;)